OpenStudy (anonymous):

1) A drive-in theatre charges $2 per vehicle plus $1 for every person in the vehicle. Find the charge for a bike with two people, by forming the equation. 2) Using the variable "a" to represent a number, write an inequality for "two times a number decreased by four that is at least five times a number".

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

is a bike considered a vehicle|||:::>???

OpenStudy (anonymous):

in this question it is i guess lol

OpenStudy (hunus):

How much do you think it would be?

OpenStudy (anonymous):

four dollars

OpenStudy (hunus):

Yup, do you know how to write the equation?

OpenStudy (anonymous):

not really no

OpenStudy (hunus):

Well for each vehicle the $2 charge doesn't change. So we know that the price P for a group of people is going to be P = 2 + ? If we have a car with one person, P would be three dollars. If we have a car with two people, P would be four dollars and so on. Do you have any idea what the '?' will be?

OpenStudy (anonymous):

2

OpenStudy (hunus):

In general, If we have x people how much will it cost for them (not including the vehicle) to get in to the theatre?

OpenStudy (anonymous):

2$ so it should be 1x right p=$2=1x

OpenStudy (anonymous):

i meant $2+1x=4

OpenStudy (hunus):

Yup :)

OpenStudy (anonymous):

will you help with second one too please

OpenStudy (hunus):

Yup

OpenStudy (anonymous):

yay thanks okay so we use variable a

OpenStudy (hunus):

Yup What do you think "two times a number decreased by four" is?

OpenStudy (anonymous):

2x-4

OpenStudy (hunus):

and what do you think "five times a number" is?

OpenStudy (anonymous):

5a

OpenStudy (anonymous):

2a-4(5a)

OpenStudy (hunus):

Yup (be sure to use 'a' on both sides) We have to account for the 'at least' statement

OpenStudy (anonymous):

did i not use a on both sides?

OpenStudy (hunus):

2a-4 '?' (5a) It will be <, <=, =, >, or >=

OpenStudy (hunus):

Was talking about when you used x in "2x-4"

OpenStudy (anonymous):

ooh thats the inequality part so maybe we could use >or= to 5a

OpenStudy (hunus):

Correct :)

OpenStudy (anonymous):

okay so the inequality is 2a-4<= 5a

OpenStudy (hunus):

2a-4 >= 5a

OpenStudy (anonymous):

thats what i meant lol sorry

OpenStudy (hunus):

That's it :)

OpenStudy (anonymous):

yay on to my next question then thanks again

OpenStudy (hunus):

No problem